OpenStudy (anonymous):

If y varies inversely as x and y = 12 when x = 3, find x when y = 18

OpenStudy (anonymous):

Since, y varies inversely as x hence \[\huge y \alpha \frac{1}{x} \rightarrow y = \frac{k}{x} \rightarrow y x = k\] where K is any constant. Let us find the value of constant K So at y = 12 when x = 3 12 * 3= = k i.e. k= 36 hence yx = 36 is the required eq of variation Now at y = 18, x=? Substituting y= 18 in eq of variation we find; 18* x= 36 x= 36/18 i.e. x= 2 is the required answer. @grayp